I can not figure out what formula to use. I have the answer in the back of the book but cannot figure out a formula to get me there. The question is: David split his $25,000 inheritance betwen Fiderlity Short Term Bond Fund with an annual yield of 5% and T Rowe price with an annual yield of 4%. If his total income for one year on the two investments was $1140, then how much did he invest in each fund? Please help - I have not done algebra in 40 years and these word problems are killing me. thank you.
Answer by fractalier(6550)

Okay, here's how we do this one:
Let x = the amount invested at 5%
Therefore 25000 - x = the amount invested at 4%.
Be clear in your mind how we get 25000 - x for the other amount.
Now the equation we use is this:
(.05)x + (.04)(25000 - x) = 1140
Let us multiply by 100 to clear the decimals (it makes it a bit easier)...
5x + 4(25000 - x) = 114000
5x + 100000 - 4x = 114000
x = 14000 the amount at 5%
25000 - x = 11000 the amount at 4%